# Implementation of time-dependent Hartree Fock in real space

### v2

@@ 3.4 Real-space calculations
$\epsilon=\frac{1}{N_{e}}\int\left|\rho^{\rm in}-\rho^{\rm out}\right|d^{3}r$ (14)
@@ 4.1 Speed-up of HF with ACE
Figure 1: Scaling of time per SCF iteration for different approaches, showing similar scaling for each but an improved pre-factor for HF when using ACE. Times are calculated as an average for the first 20 SCF iterations for the nitrosomethane molecule. Data was obtained for a fixed number of orbitals by varying the grid spacing by $0.01$ from $0.20$ to $0.09$ on a simulation domain constructed from a union of spheres of radii $5$ around each atom.
@@ 4.2 Starting point for SCF convergence of HF
The ground-state SCF calculation can be difficult to converge for HF. The default initial guess in Octopus is a linear combination of atomic orbitals (LCAO) for the wavefunctions and a superposition of atomic densities for the density. One way to accelerate this convergence is to first calculate the ground-state using LDA (faster and more easily converged) and then use that a starting point instead of using LCAO calculated in HF as the initial guess for the HF calculation, assuming that the convergence density will be similar. This approach is the standard recommendation in Octopus for hybrid functionals. To assess the convergence behaviour, we examine the relative density error as defined in equation 14.
@@
Figure 2: Convergence of HF ground-state for thioformaldehyde ($h=0.22$, $R=5$) from different starting points, measured by the relative density error defined in equation 14. First an LDA calculation (black) was performed starting from a Linear Combination of Atomic Orbitals (LCAO) and the converged results were used as initial guess for HF calculation (red). The blue curve shows convergence when the HF calculation was performed directly from LCAO.
@@ 4.3 Mixing for SCF convergence of HF
Figure 4: Convergence, as measured by the relative density error defined in equation 14, while mixing states (red and black with different mixing parameters) using linear scheme for the nitrosomethane molecule ($h=0.16$, $R=5$). Blue and magenta curves are linear mixing of potential and density, respectively, shown as a comparison to states mixing.
@@ 4.5 Benchmarking TDHF
For benchmarking purposes, we’ve chosen to compare our real-space excitation energy calculations in Octopus to those from [49] which uses the GTO code Psi4 [50]. Out of the $17$ molecules from the GTO dataset studied there (from [5]), we excluded formaldehyde, methamine and streptocyanine-c1 from our comparison due to problems with convergence of unoccupied levels in real space. We have compared the excitation energies obtained from these two different codes for IPA and TDA from TDHF as well as from TDDFT (using the BLYP functional).
The unoccupied orbitals and their symmetry is sensitive to the choice of basis set. The number of virtual orbitals obtainable depends on the number of basis functions in GTOs while in real-space it depends on the number of grid-points. For example, using aug-cc-pvtz basis set on acetaldehyde gives 496 basis functions [49], meaning a total of 496 eigenstates can be calculated. But in real-space we are not limited to that number and finer grids give us a larger number of eigenstates. The extra states obtained in real-space calculation are mostly low-lying Rydberg states. This makes the process of targeting the same excitation for comparison in GTO and real-space based approaches difficult. The indices of the states targeted in [49] are different from those of the corresponding states in a real space calculation. The size of the simulation domain also affects the nature of the calculated orbitals. Increasing the domain size results in introduction of new low-lying vacuum states which may not contribute to the absorption spectrum but nonetheless are important in understanding other kinds of response [15]. Figure 7a) shows the state targeted for the triplet excitation as obtained using GTOs, which is the fourth state after the lowest unoccupied molecular orbital (LUMO). However, in real-space, the same state index represents a Rydberg orbital as shown in figure 7b). The state resembling the GTO LUMO+4 state is found in real-space as the $30^{\rm th}$ unoccupied state. With this issue in mind, for the comparison of excitation energies against the data from [49], we have ensured that we are targeting the same excitation from both approaches by making sure the symmetry of the unoccupied orbitals matches from both approaches along with their IPA excitation energies. It is harder to determine the symmetry of the orbitals in real space. The symmetry of GTOs can be used to classify orbital symmetries, typically in the solution process. In real space, the Cartesian grid inherently breaks the system’s symmetry in most cases. In this work, we have manually examined the molecular orbitals to determine their irreducible representations.
@@
Table 2 summarizes the mean error (ME) and the mean absolute error (MAE) of RS excitation energies as compared to the GTO approach from both TDDFT and TDHF for singlets and triplets. The MAE for TDHF is 2 times larger than TDDFT for singlets and 5 times for triplets, showing that exact exchange integrals introduce a greater discrepancy. Triplets have a larger MAE than singlets, which we attribute in part to favorable cancellation of errors for the singlet between the two integrals of equation 6, whereas only one term is present for the triplet. MAEs are about an order of magnitude bigger for HF eigenvalues (IPA) than for BLYP, and HF eigenvalues’ MAE is 50-100% of the TDHF MAE. This comparison indicates that the eigenvalues, coming from the long-standing and well-tested HF implementation, rather than the TDHF integrals implemented in this work, are responsible for a large fraction of the discrepancies for TDHF. These discrepancies therefore are considered to stem from the inherent numerical differences between the RS and GTO calculations. In tables 3 and 4, we show all the triplet and singlet excitation energies, respectively, as obtained from TDDFT (TDBLYP) and TDHF, in the IPA and TDA forms, from both the GTO-based Psi4 code and the real-space (RS) Octopus code. The TBE column represents the “Theoretical Best Estimate” from GTO coupled-cluster calculations in [5].
The results are summarized in figures 8 and 9, where the error bars represent the signed deviation from the TBE. We observe that in the case of triplet excitations, TDHF with GTOs is closer to the TBE than TDHF with with real space, whereas for singlet excitations, TDHF with real space is closer to the TBE than TDHF with GTOs. These differing results for TDHF stem from differences in the accuracy of two integrals in equation 6, and cancellation of errors due to basis set and level of theory. By contrast, TDDFT results show similar deviations from the TBE for both GTO and real space.
We have also calculated the absorption spectrum using TDHF and compared it to the spectrum from TDDFT. The absorption spectra of the nitrosomethane molecule is shown in figure 10. In addition, to show that these spectra can be successfully and systematically converged, we have shown the absorption spectrum for thioformaldehyde at different grid spacings in figure 11. We see that convergence of occupied eigenvalues to  0.001 eV and unoccupied eigenvalues to  0.01 eV (figure 5) leads to convergence of singlet excitation energies within  0.1 eV: the integrals in equation 6 are more demanding to converge than the eigenvalues.
@@
Figure 8: Comparison of BLYP excitation energies, from the independent particle approximation (IPA) and TDDFT in the Tamm-Dancoff approximation (TDA), as calculated with GTOs [49] and real-space (RS) for the set of 14 small molecules. The green line represents the line $\omega^{\rm RS}=\omega^{\rm GTO}$, and error bars represent the signed deviation from the “Theoretical Best Estimate” (TBE) [5].
@@
Figure 9: Comparison of TDHF excitation energies, from the independent particle approximation (IPA) and the Tamm-Dancoff approximation (TDA), as calculated with GTOs [49] and real-space (RS) for the set of 14 small molecules. The green line represents the line $\omega^{\rm RS}=\omega^{\rm GTO}$, and error bars represent the signed deviation from the “Theoretical Best Estimate” (TBE) [5].
Figure 10: Comparison of absorption spectra of nitrosomethane molecule from TDDFT using BLYP, and TDHF, both using the TDA ($R=10$, $h=0.10$ for both TDDFT and TDHF). Lorentzian of width $0.0136$ eV used for broadening.
@@ 5 Conclusion
In this work, we have demonstrated a new implementation of linear-response TDHF in the real-space Octopus code, working in the Tamm-Dancoff approximation. For a benchmark set of small molecules, we find good agreement with results from GTOs. The discrepancy between RS and GTOs for TDHF is significantly larger than for TD-BLYP calculations, showing the more sensitive nature of the TDHF integrals. It is also larger for triplets than for singlets. The discrepancy is largely accounted for by discrepancies in the HF eigenvalues, confirming that our TDHF implementation is comparable to the pre-existing HF implementation in Octopus. These discrepancies between RS and GTOs are attributed to the unavoidable use of pseudopotentials in real space; the fact that these pseudopotentials for TDHF are generated with LDA rather than HF; possible differences in degree of convergence of the RS and GTO calculations; and different treatment of Rydberg excitations. We note that in general Rydberg excitations are better described in real space than in GTOs, where their energies may be overestimated, as seen by the negative mean error of RS vs GTOs in many cases (Table 2). Although we have only shown results from RHF here, we have also implemented TDHF for UHF and confirmed that the set of excitation obtained from RHF and UHF are identical for these closed-shell molecules. We find that convergence with respect to the real-space domain’s radius and spacing is harder than for DFT, but still achievable, for both excitation energies and absorption spectra.
@@
Since a linear-response excited-state calculation requires an accurate ground-state calculation first, we examined how to get the HF ground-state in real-space in a practical amount of time. We find that starting with a solution from DFT-LDA as a initial guess instead of LCAO directly from HF speeds up the ground-state convergence considerably. The convergence might not be possible at all in some cases without this. We also find that the best way to achieve fast convergence is by mixing states. Linear mixing of potential or density also provides a considerable increase in the rate of convergence compared to the Broyden scheme, which doesn’t seem to facilitate the convergence of the HF ground-state.
@@
The calculation of HF ground-state also requires the calculation of exact exchange which is particularly time-consuming. We have analyzed the use of ACE, an approach to speeding up the exact exchange calculation which can also be used for hybrid DFT. Using the existing implementation of ACE in Octopus, we found that for our benchmark set of molecules, the use of ACE sped up the time per SCF iteration by a factor of 7-14. We have also confirmed that the results using ACE match the results without it.
@@
This work adds TDHF to the list of features of the Octopus code, allowing users to compare various approaches in the same platform. We also note that with some further developments to the code, the ingredients from TDHF can now be used to implement hybrid functionals in TDDFT linear-response calculations [8] and other theories like ensemble DFT [51, 49].

### v3

@@ 3.4 Real-space calculations
$\epsilon=\frac{1}{N_{e}}\int\left|\rho^{\rm in}\left(\textbf{r}\right)-\rho^{\rm out}\left(\textbf{r}\right)\right|d^{3}\textbf{r}$ (15)
@@
The workflow of TDHF in Octopus is very similar to that of a standard linear-response TDDFT calculation [55]. With our implementation, we can now use flag CalculationMode = casida along with TheoryLevel = hartree_fock. For the TDA calculations shown in this study, an additional flag, CasidaTheoryLevel = tamm_dancoff, has been used. The use of ACE in the ground state and unoccupied orbitals calculations can be turned on via flag AdaptivelyCompressedExchange = yes.
@@ 4.1 Speed-up of HF with ACE
Figure 1: Scaling of time per SCF iteration for different approaches, showing similar scaling for each but an improved pre-factor for HF when using ACE. Times are calculated as an average for the first 20 SCF iterations for the nitrosomethane molecule. Data was obtained for a fixed number of orbitals by varying the grid spacing ($h$) by $0.01$ from $0.20$ to $0.09$ on a simulation domain constructed from a union of spheres of radii $5$ around each atom.
@@ 4.2 Starting point for SCF convergence of HF
The ground-state SCF calculation can be difficult to converge for HF. The default initial guess in Octopus is a linear combination of atomic orbitals (LCAO) for the wavefunctions and a superposition of atomic densities for the density. One way to accelerate this convergence is to first calculate the ground state using LDA (faster and more easily converged) and then use that a starting point instead of using LCAO calculated in HF as the initial guess for the HF calculation, assuming that the converged density will be similar. This approach is the standard recommendation in Octopus for hybrid functionals. To assess the convergence behaviour, we examine the relative density error as defined in equation 15.
@@
Figure 2: Convergence of HF ground state for thioformaldehyde ($h=0.22$, $R=5$) from different starting points, measured by the relative density error defined in equation 15. First an LDA calculation (black) was performed starting from a Linear Combination of Atomic Orbitals (LCAO) guess and the converged results were used as initial guess for HF calculation (red). The blue curve shows convergence when the HF calculation was performed directly from LCAO.
@@ 4.3 Mixing for SCF convergence of HF
Figure 4: Convergence, as measured by the relative density error defined in equation 15, while mixing states (red and black with different mixing parameters) using the linear scheme for the nitrosomethane molecule ($h=0.16$, $R=5$). Blue and magenta curves are linear mixing of potential and density, respectively, shown as a comparison to states mixing.
@@ 4.5 Benchmarking TDHF
For benchmarking purposes, we’ve chosen to compare our real-space excitation energy calculations in Octopus to those from [56] which uses the GTO code Psi4 [57]. Out of the $17$ molecules from the GTO dataset studied there (ultimately from [5], including geometries), we excluded formaldehyde, methamine and streptocyanine-c1 from our comparison due to problems with convergence of unoccupied levels in real space. We have compared the excitation energies obtained from these two different codes for IPA and TDA from TDHF as well as from TDDFT (using the BLYP functional).
The unoccupied orbitals and their symmetry is sensitive to the choice of basis set. The number of virtual orbitals obtainable depends on the number of basis functions in GTOs while in real-space it depends on the number of grid-points. For example, using the aug-cc-pvtz basis set for acetaldehyde gives 496 basis functions [56], meaning a total of 496 eigenstates can be calculated. But in real space we are not limited to that number and finer grids give us a larger number of eigenstates. The extra states obtained in real-space calculations are mostly low-lying Rydberg states. This makes the process of targeting the same excitation for comparison in GTO and real-space-based approaches difficult. The indices of the states targeted in [56] are different from those of the corresponding states in a real-space calculation. The size of the simulation domain also affects the nature of the calculated orbitals. Increasing the domain size results in introduction of new low-lying vacuum states which may not contribute to the absorption spectrum but nonetheless are important in understanding other kinds of response [16]. Figure 7a) shows the state targeted for the triplet excitation as obtained using GTOs, which is the fourth state after the lowest unoccupied molecular orbital (LUMO). However, in real space, the same state index represents a Rydberg orbital as shown in figure 7b). The state resembling the GTO LUMO+4 state is found in real space as the $30^{\rm th}$ unoccupied state. With this issue in mind, for the comparison of excitation energies against the data from [56], we have ensured that we are targeting the same excitation from both approaches by making sure the symmetry of the unoccupied orbitals matches from both approaches along with their IPA excitation energies. It is harder to determine the symmetry of the orbitals in real space. The symmetry of GTOs can be used to classify orbital symmetries, typically in the solution process. In real space, the Cartesian grid inherently breaks the system’s symmetry in most cases. In this work, we have manually examined the molecular orbitals to determine their irreducible representations.
@@
Table 2 summarizes the mean error (ME) and the mean absolute error (MAE) of RS excitation energies as compared to the GTO approach from both TDDFT and TDHF for singlets and triplets. The MAE for TDHF is 2 times larger than TDDFT for singlets and 5 times for triplets, showing that exact-exchange integrals introduce a greater discrepancy. Triplets have a larger MAE than singlets, which we attribute in part to favorable cancellation of errors for the singlet between the two integrals of equation 7, whereas only one term is present for the triplet. MAEs are about an order of magnitude bigger for HF eigenvalues (IPA) than for BLYP, and HF eigenvalues’ MAE is 50-100% of the TDHF MAE. This comparison indicates that the eigenvalues, coming from the long-standing and well-tested HF implementation, rather than the TDHF integrals implemented in this work, are responsible for a large fraction of the discrepancies for TDHF. These discrepancies therefore are considered to stem from the inherent numerical differences between the RS and GTO calculations. In tables 3 and 4, we show all the triplet and singlet excitation energies, respectively, as obtained from TDDFT (TDBLYP) and TDHF, in the IPA and TDA forms, from both the GTO-based Psi4 code and the real-space (RS) Octopus code. The TBE column represents the “Theoretical Best Estimate” from GTO coupled-cluster calculations in [5].
The results are summarized in figures 8 and 9, where the error bars represent the signed deviation from the TBE. We observe that in the case of triplet excitations, TDHF with GTOs is closer to the TBE than TDHF with real space, whereas for singlet excitations, TDHF with real space is closer to the TBE than TDHF with GTOs. These differing results for TDHF stem from differences in the accuracy of two integrals in equation 7, and cancellation of errors due to basis set and level of theory. By contrast, TDDFT results show similar deviations from the TBE for both GTO and real space.
We have also calculated the absorption spectrum using TDHF and compared it to the spectrum from TDDFT. The absorption spectra of the nitrosomethane molecule is shown in figure 10. In addition, to show that these spectra can be successfully and systematically converged, we have shown the absorption spectrum for thioformaldehyde at different grid spacings in figure 11. We see that convergence of occupied eigenvalues to $\sim 0.001$ eV and unoccupied eigenvalues to $\sim 0.01$ eV (figure 5) leads to convergence of singlet excitation energies within $\sim 0.1$ eV: the integrals in equation 7 are more demanding to converge than the eigenvalues. We also demonstrate convergence with respect to the simulation sphere $R$ in figure 12, for the absorption spectrum of diazomethane. The low-lying excitation around 1 eV is fairly well converged by $R=10$, but the higher excitations around 7 eV and above need a yet larger sphere because of their Rydberg nature and interaction with the continuum states of the vacuum.
@@
Figure 8: Comparison of BLYP excitation energies, from the independent particle approximation (IPA) and TDDFT in the Tamm-Dancoff approximation (TDA), as calculated with GTOs [56] and real space (RS) for the set of 14 small molecules. The green line represents the line $\omega^{\rm RS}=\omega^{\rm GTO}$, and error bars represent the signed deviation from the “Theoretical Best Estimate” (TBE) [5].
@@
Figure 9: Comparison of TDHF excitation energies, from the independent particle approximation (IPA) and the Tamm-Dancoff approximation (TDA), as calculated with GTOs [56] and real space (RS) for the set of 14 small molecules. The green line represents the line $\omega^{\rm RS}=\omega^{\rm GTO}$, and error bars represent the signed deviation from the “Theoretical Best Estimate” (TBE) [5].
Figure 10: Comparison of absorption spectra of nitrosomethane molecule from TDDFT using BLYP, and TDHF, both using the TDA and $R=10$, $h=0.10$. Lorentzian of width $0.0136$ eV used for broadening.
@@
Figure 12: Convergence of absorption spectra of diazomethane molecule with respect to the radius of the simulation sphere obtained from TDHF using TDA, with grid spacing $h=0.10$. Lorentzian of width $0.0136$ eV used for broadening.
@@ 5 Conclusion
In this work, we have demonstrated a new implementation of linear-response TDHF in the real-space Octopus code, working in the Tamm-Dancoff approximation. For a benchmark set of small molecules, we find good agreement with results from GTOs. The discrepancy between RS and GTOs for TDHF is significantly larger than for TD-BLYP calculations, showing the more sensitive nature of the TDHF integrals. It is also larger for triplets than for singlets. The discrepancy is largely accounted for by discrepancies in the HF eigenvalues, confirming that our TDHF implementation is comparable to the pre-existing HF implementation in Octopus. These discrepancies between RS and GTOs are attributed to the unavoidable use of pseudopotentials in real space; the fact that these pseudopotentials for TDHF are generated with LDA rather than HF; possible differences in degree of convergence of the RS and GTO calculations; and different treatment of Rydberg excitations. We note that in general Rydberg excitations are better described in real space than in GTOs, where their energies may be overestimated, as seen by the negative mean error of RS vs GTOs in many cases (Table 2). Although we have only shown results from RHF here, we have also implemented TDHF for UHF and confirmed that the set of excitations obtained from RHF and UHF is identical for these closed-shell molecules. Open-shell systems require a more sophisticated handling of spin states [58] and is beyond the scope of this study. We find that convergence with respect to the real-space domain’s radius and spacing is harder than for DFT, but still achievable, for both excitation energies and absorption spectra.
@@
Since a linear-response excited-state calculation requires an accurate ground-state calculation first, we examined how to get the HF ground state in real space in a practical amount of time. We find that starting with a solution from DFT-LDA as a initial guess instead of LCAO directly from HF speeds up the ground-state convergence considerably. The convergence might not be possible at all in some cases without this. We also find that the best way to achieve fast convergence is by mixing states. Linear mixing of potential or density also provides a considerable increase in the rate of convergence compared to the Broyden scheme, which doesn’t seem to facilitate the convergence of the HF ground state.
@@
The calculation of HF ground state also requires the calculation of exact exchange which is particularly time-consuming. We have analyzed the use of ACE, an approach to speeding up the exact-exchange calculation which can also be used for hybrid DFT. Using the existing implementation of ACE in Octopus, we found that for our benchmark set of molecules, the use of ACE sped up the time per SCF iteration by a factor of 7-14. We have also confirmed that the results using ACE match the results without it.
The parallel performance of our TDHF implementation relies on the parallel performance of the Poisson solver [33] for exchange matrix elements, using Octopus’ efficient parallelization over real-space domains [18]. TDHF can be run also with a level of parallelization over response matrix elements (equations 4, 5, 7) like in the Casida approach to TDDFT [19].
@@
This work adds TDHF to the list of features of the Octopus code, allowing users to compare various approaches in the same platform. This functionality will be available in a forthcoming release. We also note that with some further developments to the code, the ingredients from TDHF can now be used to implement hybrid functionals in TDDFT linear-response calculations [8] and other theories like ensemble DFT [59, 56].